OpenStudy (anonymous):

Solve the proportion: 16 over 17 = x over 40

OpenStudy (campbell_st):

so its \[\frac{16}{17} = \frac{x}{40}\] multiply both sides of the equation by 40 then evalaute

OpenStudy (anonymous):

i don't know how to do that lol

OpenStudy (campbell_st):

wow thats a bit sad \[\frac{16}{17} \times 40 = x\] multiplying both sides of the equation by 40 removes the denominator of 40 on the right hand side

OpenStudy (anonymous):

640/17

OpenStudy (anonymous):

i know that already but this is what i the answers A. 42.5 B. 37.6 C. -42.5 D. -37.6

OpenStudy (akashdeepdeb):

Calculate it!

OpenStudy (goformit100):

Just use the unitary method

OpenStudy (campbell_st):

what is \[640 \div 17 = ?\]

OpenStudy (anonymous):

37.64 thanks :)